# Joint Transmit and Pinching Beamforming for Pinching Antenna Systems (PASS): Optimization-Based or Learning-Based?

### v2

@@ III-B The Proposed MM-PDD Algorithm
Given $a,\theta\in\mathbb{R}$ and $c\in\mathbb{C}$, function $-\mathrm{Re}\left\{ce^{i\left(a\theta\right)}\right\}$ has Lipschitz-continuous gradient w.r.t. $\theta$, and the Lipschitz constant is given by $\varrho^{\theta}=a^{2}|c|.$
@@
From the above Lemma, the Lipschitz gradient surrogate of $L_{k,n,l}^{\mathrm{EXP}}\left(\theta_{k,n,l}\right)$ can be constructed by (1), which is defined in the bottom of this page. Furthermore, the corresponding Lipschitz constant can be given by
$\varrho_{k,n,l}^{\theta}=\phi\left|\lambda_{k,n,l}^{u}+\frac{u_{k,n,l}s_{k,n,l}}{\rho}\right|.$ (1)
@@
7:        Update $\mathbf{x}$ by solving (III-B) while fixing remaining variables, and then update $\mathbf{U}$ by (6).
8:        Update $\left\{\bm{\theta},\mathbf{S}\right\}$ by (8) using Lipschitz gradient surrogate (1a).
@@
10:     if $\left\|\mathbf{B}^{(t)}\right\|_{\infty}\leqslant 0.9\left\|\mathbf{B}^{(t-1)}\right\|_{\infty}$ then
11:        Update dual variables $\bm{\lambda}^{(t+1)}\leftarrow\bm{\lambda}^{(t)}+\frac{1}{\rho^{(t)}}\mathbf{B}^{(t)}$.
@@
13:        Update penalty parameter $\rho^{(t+1)}\leftarrow\rho^{(t)}\cdot 0.85$.
@@
15:     Algorithm terminates if $\left\|\mathbf{B}^{(t)}\right\|_{\infty}\leqslant\epsilon$.
@@
16:  $\mathbf{X},\mathbf{D}$.
$\min_{\mathbf{Q},\mathbf{D}}~{}$ $\sum_{k=1}^{K}\alpha_{k}e_{k}\left(\mathbf{Q},v_{k}^{opt}\right)\!+\!\frac{1}{2\rho}\left\|\mathbf{Q}\!\!-\!\!\mathbf{U}^{T}\bm{\Sigma}\mathbf{D}\!+\!\rho\bm{\lambda}^{q}\right\|^{2}\!\!,$ (2a)
@@
The optimal solution of $q_{k,k^{\prime}}$ can be obtained by
$q_{k,k^{\prime}}^{*}=\frac{2v_{k}\delta_{k,k^{\prime}}-\lambda_{k,k^{\prime}}^{q}+\frac{1}{\rho}\mathbf{u}_{k}^{T}\bm{\Sigma}\mathbf{d}_{k}}{2|v_{k}|^{2}+\frac{1}{\rho}},~{}\forall k,k^{\prime}\in\mathcal{K},$ (3)
[…]
$\min_{\mathbf{U}}\!\sum_{k\in\mathcal{K}}\!\left(\left\|\mathbf{R}_{k}\mathbf{u}_{k}\!+\!\bm{\zeta}_{k}\right\|^{2}\!+\!\left\|\mathbf{q}_{k}\!+\!\rho\bm{\lambda}_{k}^{q}\!-\!\mathbf{u}_{k}^{T}\bm{\Sigma}\mathbf{D}\right\|^{2}\!\right)\!,$ (5)
where $\mathbf{R}_{k}\!\!=\!\!\text{diag}\big{(}r_{k,1,1},r_{k,1,2},\dots,r_{k,N,L}\big{)}\!\in\!\mathbb{R}^{M\times M}$, and $r_{k,n,l}=r\left(x_{n,l},\bm{\eta}_{k}^{\mathrm{U}}\right)$. Moreover, $\zeta_{k,n,l}=\rho\lambda_{k,n,l}^{u}-\phi e^{-i\theta_{k,n,l}}$, and vector $\bm{\zeta}_{k}=\left[\zeta_{k,1,1},\zeta_{k,1,2},\dots,\zeta_{k,N,L}\right]^{T}\in\mathbb{C}^{M\times 1}$. Hence, the closed-form solution of $\mathbf{u}_{k}$, $\forall k\in\mathcal{K}$, can be derived as
(iv) Subproblem w.r.t. $\left\{\bm{\theta},\mathbf{S}\right\}$: The signal phases $\bm{\theta}$ and the PA-user distance $\mathbf{S}$ can be updated by solving the unconstrained convex optimization problem:
$\min_{\bm{\theta},\mathbf{S}}~{}\sum_{n,l}\sum_{k}\Bigg{[}\frac{1}{2\rho}\!\left(\theta_{k,n,l}\!-\!\kappa\!\left(s_{k,n,l}\!+\!n_{\mathrm{eff}}x_{n,l}\right)+\rho\lambda_{k,n,l}^{\theta}\right)^{2}\\ +\!\!\nabla_{\!\theta_{m,\!k}}\!L_{k,n,l}^{\mathrm{EXP}}\!\left(\!\theta_{k,n,l}^{(t-1)}\right)\theta_{k,n,l}\!+\!\frac{\varrho_{k,n,l}^{\theta}}{2}\!\left(\theta_{k,n,l}\!\!-\!\!\theta_{k,n,l}^{(t-1)}\right)^{2}\!\\ +\frac{1}{2\rho}\left(\sqrt{\left(x_{n,l}\!-\!x_{k}^{\mathrm{U}}\right)^{2}\!+\!\psi_{k,n,l}^{2}}\!-\!s_{k,n,l}+\rho\lambda_{k,n,l}^{S}\right)^{2}\!\Bigg{]}\!,$ (7)
where the optimal solution can be given by
$\theta_{k,n,l}^{*}\!\!=\!\!\frac{\varrho_{k,n,l}^{\theta}\theta_{k,n,l}^{(t-1)}\!\!-\!\!\lambda_{k,n,l}^{\theta}\!\!+\!\!\frac{\kappa}{\rho}\!\left(s_{k,n,l}\!+\!n_{\mathrm{eff}}x_{n,l}\!\right)\!\!-\!\!\nabla_{\!\theta}L^{\mathrm{EXP}}\!\!\left(\!\theta_{k,n,l}^{(t-1)}\!\right)\!}{\varrho_{k,n,l}^{\theta}\!+\!\frac{1}{\rho}}\!,$
$s_{k,n,l}^{*}\!=\!\frac{r\!\left(x_{n,l},\bm{\eta}_{k}^{\mathrm{U}}\right)\!+\!\!\rho\lambda_{k,n,l}^{S}\!+\!\!\kappa\!\left(\theta_{k,n,l}\!+\!\rho\lambda_{k,n,l}^{\theta}\!\!\!-\!\!\kappa n_{\mathrm{eff}}x_{n,l}\right)}{1\!+\!\kappa^{2}}.$ (8)
Algorithm 1 summarizes the iterative procedure of MM-PDD. By alternatively optimizing the coupled variables and address the nonconvex components using the CCCP and Lipschitz gradient surrogate, the objective value (i.e., the AL function) will be non-increasing through each BCD update in the inner loop [16, 17]. Therefore, MM-PDD is guaranteed to converge to a stationary point according to the analysis in [19]. However, obtaining a high quality solution $\left\{\mathbf{X}^{*},\mathbf{D}^{*}\right\}$ in this way is very time consuming, and is prone to getting stuck in local optimums, which motivates us to further develop the learning-based approach.
@@ IV-B1 Fundamentals of KDL
$\bm{\varphi}\!\left(\mathbf{D},\bm{\tau},\bm{\lambda},\bm{\mu}\right)\!\!\triangleq\!\!\left[\!\begin{array}[]{c}\nabla_{\mathbf{D}}F_{0}\!\left(\mathbf{D};\mathbf{X},\mathbf{z}\right)\!+\!\mathbf{J}_{\mathbf{f},\mathbf{D}}^{T}\bm{\lambda}\\ \mathrm{diag}\left(\bm{\lambda}\right)\mathbf{f}\left(\mathbf{D}\right)\end{array}\!\right]\!=\!\left[\!\begin{array}[]{c}\mathbf{0}\\ \mathbf{0}\end{array}\!\right],$ (14)
where $\mathbf{J}_{\mathbf{f},\mathbf{D}}^{T}=\left[\nabla_{\mathbf{D}}f_{1}\left(\mathbf{D}\right),\nabla_{\mathbf{D}}f_{2}\left(\mathbf{D}\right),\dots,\nabla_{\mathbf{D}}f_{F_{D}}\left(\mathbf{D}\right)\right]$ is the Jacobian matrix of constraint functions $\mathbf{f}\left(\mathbf{D};\mathbf{z}\right)$ over $\mathbf{D}$.
@@ IV-C The Proposed KDL-Transformer Algorithm
Following existing LLM structure, the developed Transformer consists of a $J_{\mathrm{E}}$-layer encoder $\mathcal{E}$ and a $J_{\mathrm{D}}$-layer decoder $\mathcal{D}$. Compared to vanilla Transformer that exploits masked attention the the decoder, we tailor the Tranformer structure to utilize bidirectional attention mechnisms, thus modelling global dependence in both the encoder and the decoder: (i) Both the encoder and decoder adopt bidirectional self-attention mechanisms, which characterizes the inter-user dependence in the input sequence and the inter-PA dependence in the output sequence, respectively. (ii) Moreover, the cross-attention mechanism is adopted between the encoder and the decoder to learn the CSI-beamforming dependence, thus refining the transmit bemaforming and pinching beamforming coefficients in correspondence to different CSI. The detailed designs of encoder and decoder are illustrated in the sequel.
@@ IV-C1 Encoder Design
The self-attention mechanism of Transformer allows each element (which is also known as token) in the sequence to interact with other elements to model their global dependencies. To calculate the attention weights for the vectorized CSI input $\mathbf{z}=\left[\mathbf{x}^{\mathrm{U}},\mathbf{y}^{\mathrm{U}}\right]\in\mathbb{R}^{2K\times 1}$, we need to first embed $\mathbf{z}$ into a high-dimension matrix $\mathbf{Z}_{\mathrm{enc}}\in\mathbb{R}^{2K\times N_{\mathrm{model}}}$, where $N_{\mathrm{model}}$ denotes the feature embedding dimension. This can be realized using a learnable linear transform $\mathbf{w}_{\mathrm{E}}\in\mathbb{R}^{N_{\mathrm{model}}\times 1}$ for $\mathbf{z}$:
$\mathbf{Z}_{\mathrm{enc}}=\mathbf{z}^{\top}\mathbf{w}_{\mathrm{E}}.$ (23)
@@
The encoder exploits a $J_{\mathrm{E}}$-layer structure. In each encoding layer, given the embedding feature (i.e., hidden state) $\mathbf{Z}_{\mathrm{enc}}$, Transformer generates three different matrices, named queue $\mathbf{Y}_{i}^{Q}$, key $\mathbf{Y}_{j}^{K}$, and value $\mathbf{Y}_{j}^{V}$. Specifically, queue $\mathbf{Y}_{i}^{Q}$ indicates which information requires attentions from the current elements, and key $\mathbf{Y}_{j}^{K}$ signifies what information can be provided by each element. Hence, the self-attention score $\mathbf{Y}^{Q}\left(\mathbf{Y}^{K}\right)^{\top}$ is defined as the inner product of queue and key, which denotes the correlation between elements. Moreover, value $\mathbf{Y}_{j}^{V}$ is the representation of the actual information maintained by each element, which will be multiplied with the self-attention score for hidden state update.
We employ multiple attention heads to capture different inter-dependence patterns. Each attention head $i$ individually computes queue, key, and value matrices by
$\mathbf{Y}_{i}^{K}\!=\!\mathbf{Z}_{\mathrm{enc}}\!\mathbf{W}_{i}^{K},\quad\mathbf{Y}_{i}^{Q}\!=\!\mathbf{Z}_{\mathrm{enc}}\!\mathbf{W}_{i}^{q},\quad\mathbf{Y}_{i}^{V}\!=\!\mathbf{Z}_{\mathrm{enc}}\!\mathbf{W}_{i}^{V},$ (24)
@@
where $I$ denotes the total number of attention heads, $\mathbf{W}_{i}^{Q},\mathbf{W}_{i}^{K},\mathbf{W}_{i}^{V}\in\mathbb{R}^{N_{\text{model}}\times N_{\text{model}}}$ denotes the learnable weight matrices. By normalizing the self-attention scores using the Softmax function and multiplying it with value matrix, each head $i$ obtains the attention weights as follows:
@@
$\text{head}_{i}\!\!=\!\!Attn(\mathbf{Y}_{i}^{Q},\mathbf{Y}_{i}^{K},\mathbf{Y}_{i}^{V})\!\!=\!\!\text{softmax}\!\!\left(\frac{\mathbf{Y}_{i}^{Q}\!\left(\mathbf{Y}_{i}^{K}\right)^{\top}}{\sqrt{N_{\mathrm{model}}}}\!\right)\!\mathbf{Y}_{i}^{V}\!.$ (25)
@@
Then, the encoder concatenates attention weights from different heads, and computes multi-head features through a learnable matrix $\mathbf{W}^{O}$ and a feed-forward network (FFN):
$\begin{split}&\text{MultiHead}\left(\left\{\mathbf{Y}_{i}^{Q}\right\},\left\{\mathbf{Y}_{i}^{K}\right\},\left\{\mathbf{Y}_{i}^{V}\right\}\right)=\\ &\quad\quad\mathrm{FFN}\left(\text{Concat}(\text{head}_{1},\text{head}_{2},...,\text{head}_{I})\mathbf{W}^{O}\right).\end{split}$ (26)
The hidden state of each encoding layer will be updated based on the multi-head features using a residual connection layer:
@@
$\mathbf{Z}_{\mathrm{enc}}\leftarrow\mathbf{Z}_{\mathrm{enc}}+\text{MultiHead}\left(\left\{\mathbf{Y}_{i}^{Q}\right\},\left\{\mathbf{Y}_{i}^{K}\right\},\left\{\mathbf{Y}_{i}^{V}\right\}\right),$ (27)

### v3

@@ III-B The Proposed MM-PDD Algorithm
Given $a,\theta\in\mathbb{R}$ and $c\in\mathbb{C}$, function $\mathrm{Re}\left\{ce^{i\left(a\theta\right)}\right\}$ has Lipschitz-continuous gradient w.r.t. $\theta$, and the Lipschitz constant is given by $\varrho^{\theta}=a^{2}|c|.$
@@
From the above Lemma, the Lipschitz gradient surrogate of $L_{k,n,l}^{\mathrm{ex}}\left(\theta_{k,n,l}\right)$ can be constructed by (1), which is defined in the bottom of this page. Furthermore, the corresponding Lipschitz constant can be given by
$\varrho_{k,n,l}^{\theta}=\phi\left|\lambda_{k,n,l}^{u}+\frac{u_{k,n,l}r_{k,n,l}}{\rho}\right|.$ (1)
$\min_{\mathbf{Q},\mathbf{D}}$~ $\sum_{k=1}^{K}\alpha_{k}e_{k}\left(\mathbf{Q},v_{k}^{opt}\right)\!+\!\frac{1}{2\rho}\left\|\mathbf{Q}\!\!-\!\!\mathbf{U}^{\top}\bm{\Sigma}\mathbf{D}\!+\!\rho\bm{\lambda}^{q}\right\|^{2}\!\!,$ (2a)
$\text{s.t.}\quad\text{(C3)}.$ (2b)
The optimal solution of $q_{k,k^{\prime}}$ can be obtained by
$q_{k,k^{\prime}}^{*}=\frac{2v_{k}\delta_{k,k^{\prime}}-\lambda_{k,k^{\prime}}^{q}+\frac{1}{\rho}\mathbf{u}_{k}^{\top}\bm{\Sigma}\mathbf{d}_{k}}{2|v_{k}|^{2}+\frac{1}{\rho}},~\forall k,k^{\prime}\in\mathcal{K},$ (3)
$\text{s.t.}\quad\text{(C1), (C2)}.$ (4)
$\min_{\mathbf{U}}\!\sum_{k\in\mathcal{K}}\!\left(\left\|\mathbf{R}_{k}\mathbf{u}_{k}\!+\!\bm{\zeta}_{k}\right\|^{2}\!+\!\left\|\mathbf{q}_{k}\!+\!\rho\bm{\lambda}_{k}^{q}\!-\!\mathbf{u}_{k}^{\top}\bm{\Sigma}\mathbf{D}\right\|^{2}\!\right)\!,$ (5)
where $\mathbf{R}_{k}\!\!=\!\!\text{diag}\big(r_{k,1,1},r_{k,1,2},\dots,r_{k,N,L}\big)\!\in\!\mathbb{R}^{M\times M}$, and $r_{k,n,l}=r\left(x_{n,l},\bm{\eta}_{k}^{\mathrm{U}}\right)$. Moreover, $\zeta_{k,n,l}=\rho\lambda_{k,n,l}^{u}-\phi e^{-i\theta_{k,n,l}}$, and vector $\bm{\zeta}_{k}=\left[\zeta_{k,1,1},\zeta_{k,1,2},\dots,\zeta_{k,N,L}\right]^{\top}\in\mathbb{C}^{M\times 1}$. Hence, the closed-form solution of $\mathbf{u}_{k}$, $\forall k\in\mathcal{K}$, can be derived as
(iv) Subproblem w.r.t. $\bm{\theta}$: The signal phases $\bm{\theta}$ and the PA-user distance $\mathbf{S}$ can be updated by solving the unconstrained convex optimization problem:
$\min_{\bm{\theta},\mathbf{S}}~\sum_{n,l}\sum_{k}\Bigg[\frac{1}{2\rho}\!\left(\theta_{k,n,l}\!-\!\kappa\!\left(r_{k,n,l}\!+\!n_{\mathrm{eff}}x_{n,l}\right)+\rho\lambda_{k,n,l}^{\theta}\right)^{2}\\ +\!\!\nabla_{\theta}L_{k,n,l}^{\mathrm{ex}}\!\left(\theta_{k,n,l}^{(t-1)}\right)\!\!\theta_{k,n,l}\!+\!\frac{\varrho_{k,n,l}^{\theta}}{2}\!\left(\theta_{k,n,l}\!-\!\theta_{k,n,l}^{(t-1)}\right)^{2}\!\Bigg]\!,$ (7)
where the optimal solution can be given by
$\theta_{k,n,l}^{\star}\!\!=\!\!\frac{\varrho_{k,n,l}^{\theta}\theta_{k,n,l}^{(t-1)}\!\!-\!\!\lambda_{k,n,l}^{\theta}\!\!+\!\!\frac{\kappa}{\rho}\!\left(r_{k,n,l}\!+\!n_{\mathrm{eff}}x_{n,l}\!\right)\!\!-\!\!\nabla_{\!\!\theta}L^{\!\mathrm{ex}}\!\!\left(\!\theta_{k,n,l}^{(t-1)}\!\right)\!}{\varrho_{k,n,l}^{\theta}\!+\!\frac{1}{\rho}}\!.$ (8)
@@
7:   Update $\mathbf{x}$ by (III-B) while fixing remaining variables.
8:   Update $\mathbf{U}$ by (6), and then update $\bm{\theta}$ by (8) via Lipschitz gradient surrogate (1a).
@@
10:  if $\left\|\mathbf{B}^{(i)}\right\|_{\infty}\leqslant\varepsilon^{(i)}\!\triangleq\!\gamma\left\|\mathbf{B}^{(i-1)}\right\|_{\infty}$ then
11:   Update $\bm{\lambda}^{(i+1)}\leftarrow\bm{\lambda}^{(i)}+\frac{1}{\rho^{(i)}}\mathbf{B}^{(i)}$, $\rho^{(i+1)}\leftarrow\rho^{(i)}$.
@@
13:   Update $\rho^{(i+1)}\leftarrow\varsigma\rho^{(i)}$, $\bm{\lambda}^{(i+1)}\leftarrow\bm{\lambda}^{(i)}$.
@@
15:  Algorithm terminates if $\left\|\mathbf{B}^{(i)}\right\|_{\infty}\leqslant\varepsilon$.
@@
III-C Algorithm Design and Theoretical Analysis
Algorithm 1 summarizes the iterative procedure of MM-PDD [^1] In each PDD iteration $i$, the residual tolerance $\varepsilon^{(i)}=\gamma\|\mathbf{B}^{(i)}\|_{\infty}$ controls whether to update the dual variables or to decrease the penalty parameter (by $\rho^{(i+1)}\leftarrow\varsigma\rho^{(i)}$), where $0<\gamma,\varsigma<1$ enforces equality residuals $\|\mathbf{B}^{i}\|_{\infty}\to 0$ as $i$ grows. We empirically set $\gamma=0.9$ and $\varsigma=0.85$, which achieves an acceptable convergence speed without sacrificing the solution quality. .
The computational complexity of the proposed MM-PDD algorithm can be analyzed as follows. In each inner iteration, the updates of $\{v_{k},\alpha_{k}\}$ require evaluating the effective channels $\mathbf{h}_{k}(\mathbf{X})\mathbf{G}(\mathbf{X})$, $\forall k$, which scales linearly with the number of pinching antennas as $\mathcal{O}(KM)$. The update of $\{\mathbf{D},\mathbf{Q}\}$ involves solving a convex quadratic subproblem in $\mathbf{D}$, whose time complexity using interior-point methods is given by $\mathcal{O}(N^{3.5}K^{3.5})$, and updating $\mathbf{Q}$ incurs a time complexity $\mathcal{O}(KM+K^{2})$. The update of the pinching deployment $\mathbf{X}$ relies on solving a convex quadratic program with $NL$ variables, which typically dominates the iteration cost with $\mathcal{O}(N^{3.5}L^{3.5})$. The update of $\mathbf{U}$ requires $K$ independent $M\times M$ linear system inversions, with overall computational complexity $O(KM^{3})$. Finally, the updates of $\bm{\theta}$ is closed-form and only costs $\mathcal{O}(KM)$. Therefore, the time complexity of MM-PDD with $N_{\mathrm{it}}$ inner iterations can be given by $\mathcal{O}\big(N_{\mathrm{it}}\big((K^{3.5}+L^{3.5})N^{3.5}+M^{3}K+K^{2}+2KM\big)\big)$.
The developed MM-PDD is mathematically guaranteed to converge to a stationary solution, as proven below.
Lemma 3:
Algorithm 1 converges to a stationary solution (i.e., a KKT point) of problem (P0).
Proof.
See Appendix -C. ∎
Despite the strict theoretical guarantees, obtaining a high quality solution $\left\{\mathbf{X}^{*},\mathbf{D}^{*}\right\}$ by MM-PDD is highly complex, and is prone to getting stuck in local optima. This motivates us to further develop the learning-based approach.
@@ IV-B1 Fundamentals of KDL
$\bm{\varphi}\!\left(\mathbf{D},\bm{\tau},\bm{\lambda},\bm{\mu}\right)\!\!\triangleq\!\!\left[\!\begin{array}[]{c}\nabla_{\mathbf{D}}F_{0}\!\left(\mathbf{D};\mathbf{X},\mathbf{z}\right)\!+\!\mathbf{J}_{\mathbf{f},\mathbf{D}}^{\top}\bm{\lambda}\\ \mathrm{diag}\left(\bm{\lambda}\right)\mathbf{f}\left(\mathbf{D}\right)\end{array}\!\right]\!=\!\left[\!\begin{array}[]{c}\mathbf{0}\\ \mathbf{0}\end{array}\!\right],$ (14)
where $\mathbf{J}_{\mathbf{f},\mathbf{D}}^{\top}=\left[\nabla_{\mathbf{D}}f_{1}\left(\mathbf{D}\right),\nabla_{\mathbf{D}}f_{2}\left(\mathbf{D}\right),\dots,\nabla_{\mathbf{D}}f_{F_{D}}\left(\mathbf{D}\right)\right]$ is the Jacobian matrix of constraint functions $\mathbf{f}\left(\mathbf{D};\mathbf{z}\right)$ over $\mathbf{D}$.
@@ IV-C The Proposed KDL-Transformer Algorithm
Figure 2: Diagram of the proposed KDL-Transformer.
Following existing LLM structure, the developed Transformer consists of a $J_{\mathrm{E}}$-layer encoder $\mathcal{E}$ and a $J_{\mathrm{D}}$-layer decoder $\mathcal{D}$, as shown in Fig. 2. The Transformer structure is tailored to utilize bidirectional attention mechanisms, thus modelling global dependence in both the encoder and the decoder: (i) Both the encoder and decoder adopt bidirectional self-attention mechanisms, which characterizes the inter-user dependence in the input sequence and the inter-PA dependence in the output sequence, respectively. (ii) Moreover, the cross-attention mechanism is adopted between the encoder and the decoder to learn the CSI-beamforming dependence, thus refining the transmit bemaforming and pinching beamforming coefficients in correspondence to CSI conditions. The detailed designs can be illustrated in the sequel.
@@ IV-C1 Encoder Design
The self-attention mechanism of Transformer allows each element (which is also known as token) in the sequence to interact with other elements to model their global dependencies. Given the vectorized CSI input $\mathbf{z}\!=\!\big\{\mathbf{x}^{\mathrm{U}},\mathbf{y}^{\mathrm{U}}\big\}\!\in\!\mathbb{R}^{2K\times 1}$, we first embed $\mathbf{z}$ into feature $\mathbf{Z}^{\mathrm{E}}\in\mathbb{R}^{2K\times N_{\mathrm{model}}}$ by sinusoidal position embedding:
$\mathbf{Z}^{\mathrm{E}}=\mathbf{w}^{\mathrm{E}}\mathbf{z}^{\top}+\mathbf{PE},$ (23)
where $N_{\mathrm{model}}$ denotes the feature embedding dimension, and $\mathbf{w}^{\mathrm{E}}\in\mathbb{R}^{N_{\mathrm{model}}\times 1}$ is a learnable linear transform. $\mathbf{PE}$ represents the fixed sinusoidal position embedding that retains the sequential order of each token, which is defined as $\mathbf{PE}(i,2j)=\sin\left(\frac{i}{10000^{2j/N{\mathrm{model}}}}\right)$, $\mathbf{PE}(i,2j+1)=\cos\left(\frac{i}{10000^{2j/N{\mathrm{model}}}}\right)$, $j=0,1,\dots,N_{\mathrm{model}}/2-1$.
@@
The encoder exploits a $J_{\mathrm{E}}$-layer structure. In each encoding layer, given the embedding feature (i.e., hidden state) $\mathbf{Z}^{\mathrm{E}}$, Transformer generates three different matrices, named query $\mathbf{Y}_{i}^{\mathrm{Q}}$, key $\mathbf{Y}_{j}^{\mathrm{K}}$, and value $\mathbf{Y}_{j}^{\mathrm{V}}$. Specifically, query $\mathbf{Y}_{i}^{\mathrm{Q}}$ indicates which information requires attentions by each element, and key $\mathbf{Y}_{j}^{\mathrm{K}}$ signifies what information can be provided by each element. Hence, the self-attention score $\mathbf{Y}^{\mathrm{Q}}\left(\mathbf{Y}^{\mathrm{K}}\right)^{\top}$, defined as the inner product of query and key, denotes the correlation between elements. Moreover, value $\mathbf{Y}_{j}^{\mathrm{V}}$ represents the actual information maintained by each element, which will be multiplied with the self-attention score for hidden state update. We employ multiple attention heads to capture different inter-dependence patterns. Each attention head $i$ individually computes query, key, and value matrices by
$\mathbf{Y}_{i}^{\mathrm{K}}\!=\!\mathbf{Z}^{\mathrm{E}}\mathbf{W}_{i}^{\mathrm{K}},\quad\mathbf{Y}_{i}^{\mathrm{Q}}\!=\!\mathbf{Z}^{\mathrm{E}}\mathbf{W}_{i}^{\mathrm{Q}},\quad\mathbf{Y}_{i}^{\mathrm{V}}\!=\!\mathbf{Z}^{\mathrm{E}}\mathbf{W}_{i}^{\mathrm{V}},$ (24)
@@
where $I$ denotes the total number of attention heads, $\mathbf{W}_{i}^{Q},\mathbf{W}_{i}^{K},\mathbf{W}_{i}^{V}\in\mathbb{R}^{N_{\text{model}}\times N_{\text{model}}}$ denotes the learnable weight matrices. By normalizing the self-attention scores using the Softmax function and multiplying it with value matrix, each head $i$ obtains attention scores as follows:
@@
$\text{head}_{i}\!\!=\!\!Attn(\mathbf{Y}_{i}^{\mathrm{Q}},\mathbf{Y}_{i}^{\mathrm{K}},\mathbf{Y}_{i}^{\mathrm{V}})\!\!=\!\!\text{softmax}\!\!\left(\tfrac{\mathbf{Y}_{i}^{\mathrm{Q}}\!\left(\mathbf{Y}_{i}^{\mathrm{K}}\right)^{\top}}{\sqrt{N_{\mathrm{model}}}}\!\right)\!\mathbf{Y}_{i}^{\mathrm{V}}\!.$ (25)
@@
Then, the encoder concatenates attention weights from different heads, and computes multi-head features through a learnable matrix $\mathbf{W}^{O}$ and a feed-forward network (FFN) $\mathcal{W}_{\mathrm{FFN}}$. The hidden state of each encoding layer is updated by the multi-head features using a residual connection layer:
@@
$\mathbf{Z}^{\mathrm{E}}\!\leftarrow\!\mathbf{Z}^{\mathrm{E}}\!+\!\mathcal{W}_{\mathrm{FFN}}\left(\text{Concat}(\text{head}_{1},\text{head}_{2},...,\text{head}_{I})\mathbf{W}^{O}\right),$ (26)